# Exploring Multidimensional Checkworthiness: Designing AI-assisted Claim Prioritization for Human Fact-checkers

### v1

@@ 5.3. Study Limitations
The advent of LLMs has created tremendous excitements about democratizing AI capabilities. Our study empowers fact-checkers to create zero-code, custom search filters alongside a standard search box with four checkworthiness filters. Search engines have performed best for keyword-oriented queries, and users thus traditionally assume and write keyword queries. However, search engine capabilities have progressed tremendously in understanding more verbose and complex queries (Bendersky and Croft, 2008; Gupta and Bendersky, 2015), and today’s search engines increasingly incorporate the latest LLM capabilities to support more powerful query interpretation (goo, 2024). However, our prototype’s search box was only implemented semantic search using SentenceBERT (Section 3.3.2). Future work could explore the integration of LLMs directly into the search box to enhance query interpretation without additional custom filters.
@@ 6. Conclusion
With so many potentially false claims circulating online, claim prioritization is key to intelligently allocating limited human resources for fact-checking. Our study frames claim prioritization through the lens of IR: just as relevance is multidimensional, with many factors influencing which search results a user deems relevant, checkworthiness is also multi-faceted, subjective, and even personal, with many factors influencing how fact-checkers prioritize claims to check.
@@ Appendix A Claim prioritization task
Imagine that at the beginning of your everyday fact-checking, you want to find a set of claims across diverse sub-topics relevant to COVID-19. Finally, you want to recommend 3 claim candidates to editors from the set of what you already found. Specific steps are:
@@ Appendix E Preliminary findings
This section describes our preliminary findings from pilot tests to identify appropriate search tasks for fact-checkers. These tasks aim to help fact-checkers prioritize claims based on various dimensions of checkworthiness. It is well known in IR that user information needs often evolve over the course of a search session, as they encounter and explore new information that expands their initial understanding of a topic (Jiang et al., 2014). As these information needs evolve, the criteria used to determine the search relevance adjust accordingly. A classic example demonstrated by Tang and Solomon (1998) is that users seek information driven by their broad interests and curiosity during the early stage of random browsing. Then, they apply more specific, case-based relevance criteria after better understanding their search objectives. These relevance factors are more related to knowledge construction and problem-solving. In the context of claim prioritization, this suggests that the relative importance that fact-checkers ascribe to different checkworthy dimensions used in their information-seeking may similarly evolve as they search.
@@ Appendix F Implicit ranking
Section 3.4.4 mentioned that our pre-screening survey asked two related questions about the relative importance of different checkworthy dimensions. In this section, we compare how fact-checkers answers these two related questions.
First, we asked fact-checkerrs to answer a 5-point Likert scale rating question about the Perceived importance of each checkworthy dimension: “This is an important factor resulting in the final fact-checked claim.”
@@
Complementing this rating question, we further asked participants to implicitly rank the relative importance of the four dimensions in an open-ended format: “Considering these four criteria, how would you characterize their relative importance compared to one another? Please rank these criteria from most to least important.” Because these answers were open-ended, participants often mentioned only a subset of the dimensions and used free-form language requiring manual analysis. Most participants identified several dimensions as “most or equally important,” while also indicating which dimensions they considered “least important.” We counted such responses and provide these counts in Table 10.
@@ Appendix H Additional results
Verifiable 4.13 — 3.69 9.0(0.083) 4.13 — 4.31 12.0(0.365) 4.13 — 4.25 17.0(0.492)
Likely false - - 3.69 — 4.31 0.0(0.003) 3.69 — 4.25 10.5(0.033)
Likely harmful - - - - 4.31 — 4.25 14.0(0.500)
@@
Verifiable 0.77 — 0.73 24.0(0.239) 0.77 — 0.62 19.0(0.063) 0.77 — 0.39 7.0(0.002)
Likely false - - 0.73 — 0.62 27.0(0.196) 0.73 — 0.39 8.5(0.005)
Likely harmful - - - - 0.62 — 0.39 11.0(0.015)
@@
Usefulness Using this tool would enable me to accomplish task more quickly 4.13(0.52) — 4
Using this tool would improve my job performance. 4.07(0.59) — 4
Using this tool in my job would increase my productivity. 4.00(0.76) — 4
Using this tool would enhance my effectiveness on the job. 4.07(0.59) — 4
Using this tool would make it easier to do my job. 4.27(0.59) — 4
I would find this tool useful in my job. 4.40(0.51) — 4
Ease of use Learning to operate the tool would be easy for me. 4.38(0.89) — 5
I would find it easy to get this tool to do what I want it to do. 4.06(0.77) — 4
My interaction with the tool would be clear and understandable. 4.19(0.75) — 4
I would find this tool would be clear and understandable. 4.38(0.62) — 4
It would be easy for me to become skillful at using this tool. 4.25(0.77) — 4
I found this tool easy to use. 4.19(1.17) — 4

### v2

@@ 5.3. Study Limitations
The advent of LLMs has created tremendous excitement about democratizing AI capabilities. Our study empowers fact-checkers to create zero-code, custom search filters alongside a standard search box with four checkworthiness filters. Search engines have performed best for keyword-oriented queries, and users thus traditionally assume and write keyword queries. However, search engine capabilities have progressed tremendously in understanding more verbose and complex queries (Bendersky and Croft, 2008; Gupta and Bendersky, 2015), and today’s search engines increasingly incorporate the latest LLM capabilities to support more powerful query interpretation (goo, 2024). However, our prototype’s search box only implemented semantic search using SentenceBERT (Section 3.3.2). Future work could explore the integration of LLMs directly into the search box to enhance query interpretation without additional custom filters.
@@ 6. Conclusion
With so many potentially false claims circulating online, claim prioritization is key to intelligently allocating limited human resources for fact-checking. Our study perceives claim prioritization through the lens of IR: just as relevance is multidimensional, with many factors influencing which search results a user deems relevant, checkworthiness is also multi-faceted, subjective, and even personal, with many factors influencing how fact-checkers prioritize claims to check.
@@ Appendix A Claim prioritization task
Imagine that at the beginning of your everyday fact-checking, you want to find a set of claims across diverse sub-topics relevant to COVID-19. Finally, you want to recommend 3 claim candidates to editors from the list of candidates you have already found. Specific steps are:
@@ Appendix E Preliminary findings
This section describes our preliminary findings from pilot tests to identify appropriate search tasks for fact-checkers. These tasks aim to help fact-checkers prioritize claims based on various dimensions of checkworthiness. It is well known in IR that user information needs often evolve over the course of a search session as they encounter and explore new information that expands their initial understanding of a topic (Jiang et al., 2014). As this information needs to evolve, the criteria used to determine the search relevance adjust accordingly. A classic example demonstrated by Tang and Solomon (1998) is that users seek information driven by their broad interests and curiosity during the early stage of random browsing. Then, they apply more specific, case-based relevance criteria after better understanding their search objectives. These relevance factors are more related to knowledge construction and problem-solving. In the context of claim prioritization, this suggests that the relative importance that fact-checkers ascribe to different checkworthy dimensions used in their information-seeking may similarly evolve as they search.
@@ Appendix F Implicit ranking
Section 3.4.4 mentioned that our pre-screening survey asked two related questions about the relative importance of different checkworthy dimensions. In this section, we compare how fact-checkers answer these two related questions.
First, we asked fact-checkers to answer a 5-point Likert scale rating question about the Perceived importance of each checkworthy dimension: “This is an important factor resulting in the final fact-checked claim.”
@@
Complementing this rating question, we further asked participants to implicitly rank the relative importance of the four dimensions in an open-ended format: “Considering these four criteria, how would you characterize their relative importance compared to one another? Please rank these criteria from most to least important.” Because these answers were open-ended, participants often mentioned only a subset of the dimensions and used free-form language requiring manual analysis. Most participants identified several dimensions as “most or equally important,” while also indicating which dimensions they considered “least important.” We counted such responses and provided these counts in Table 10.
@@ Appendix H Additional results
Verifiable 4.13 — 3.69 9.00(0.08) 4.13 — 4.31 12.00(0.37) 4.13 — 4.25 17.00(0.49)
Likely false - - 3.69 — 4.31 0.0(0.00) 3.69 — 4.25 10.50(0.03)
Likely harmful - - - - 4.31 — 4.25 14.00(0.50)
@@
Verifiable 0.77 — 0.73 24.00(0.24) 0.77 — 0.62 19.00(0.06) 0.77 — 0.39 7.00(0.00)
Likely false - - 0.73 — 0.62 27.00(0.20) 0.73 — 0.39 8.50(0.01)
Likely harmful - - - - 0.62 — 0.39 11.00(0.02)
@@
Usefulness Using this tool would enable me to accomplish tasks more quickly 4.13(0.52) 4
Using this tool would improve my job performance. 4.07(0.59) 4
Using this tool in my job would increase my productivity. 4.00(0.76) 4
Using this tool would enhance my effectiveness on the job. 4.07(0.59) 4
Using this tool would make it easier to do my job. 4.27(0.59) 4
I would find this tool useful in my job. 4.40(0.51) 4
Ease of use Learning to operate the tool would be easy for me. 4.38(0.89) 5
I would find it easy to get this tool to do what I want it to do. 4.06(0.77) 4
My interaction with the tool would be clear and understandable. 4.19(0.75) 4
I would find this tool would be clear and understandable. 4.38(0.62) 4
It would be easy for me to become skillful at using this tool. 4.25(0.77) 4
I found this tool easy to use. 4.19(1.17) 4